Firm C currently has 250,000 shares outstanding with current market value of $42.00 per share and generates an annual EBIT of $1,250,000. Firm C also has $1 million of debt outstanding. The current cost of equity is 8 percent and the current cost of debt is 5 percent. The firm is considering issuing another $2 million of debt and using the proceeds of the debt issue to repurchase shares (a pure capital structure change). It is estimated that the cost of the new debt will be 5.5 percent and that the cost of equity will rise to 8.48 percent with the additional debt. The marginal tax rate is 30 percent.

QUESTION: What is the firm's net income after the recapitalization? Note: The total interest costs that must be subtracted from EBIT must be calculated in two parts and then added: Kdo Do + KaiD
Answer #1

The calculation is as follows:

Interest on $1 million debt (D0) = D0 * Kd0

                                           = $1,000,000 * 5%

                                          = $50,000

Interest on $2 million debt (D1) = D1* Kd1

                                                    = $2,000,000 * 5.50%

                                                    = $110,000

Total Interest = (D0 * Kd0) + (D1* Kd1)

                       = $50,000 + 110,000

                      = $160,000

Recapitalized net income = (EBIT – Total interest) (1 – tax rate)

                                         = ($1,250,000 – $160,000) * (1 - 30%)

                                        = $1,090,000 * (1 - 30%)

                                       = $763,000

Hence, the firm’s net income (recapitalized) is $763,000.